a car travelling south at 36 m/s sounds a 456 hz horn . what frequency is heard by the driver in an approaching truck traveling north at 25 m/s? what frequency is heard by the driver if he is heading south away from the car at 25 m/s?


1
Expert's answer
2021-01-14T10:39:18-0500

a) We can find the frequency which is heard by the driver in the approaching truck from the Doppler Shift formula:


"f_o=f_s\dfrac{v+v_o}{v-v_s},"

here, "f_o" is the frequency which is heard by the driver in the approaching truck, "f_s=456\ Hz" is the source frequency, "v=340\ \dfrac{m}{s}" is the speed of the sound, "v_s=36\ \dfrac{m}{s}" is the speed of the source (or the car), "v_o=25\ \dfrac{m}{s}" is the speed of the observer (or the speed of the driver in the truck).

Then, we get:


"f_o=456\ Hz\cdot\dfrac{340\ \dfrac{m}{s}+25\ \dfrac{m}{s}}{340\ \dfrac{m}{s}-36\ \dfrac{m}{s}}=547.5\ Hz."

b) We can find the frequency which is heard by the driver in the truck if he is heading south away from the car from the Doppler Shift formula:


"f_o=f_s\dfrac{v-v_o}{v-v_s},"

here, "f_o" is the frequency which is heard by the driver in the truck heading south away from the car, "f_s=456\ Hz" is the source frequency, "v=340\ \dfrac{m}{s}" is the speed of the sound, "v_s=36\ \dfrac{m}{s}" is the speed of the source (or the car), "v_o=25\ \dfrac{m}{s}" is the speed of the observer (or the speed of the driver in the truck).

Then, we get:


"f_o=456\ Hz\cdot\dfrac{340\ \dfrac{m}{s}-25\ \dfrac{m}{s}}{340\ \dfrac{m}{s}-36\ \dfrac{m}{s}}=472.5\ Hz."

Answer:

(a) "f_o=547.5\ Hz."

(b) "f_o=472.5\ Hz."